### v2

@@ Abstract
We revisit the relationship between attention mechanisms and large kernel ConvNets in visual transformers and propose a new spatial attention named Large Kernel Convolutional Attention (LKCA). It simplifies the attention operation by replacing it with a single large kernel convolution. LKCA combines the advantages of convolutional neural networks and visual transformers, possessing a large receptive field, locality, and parameter sharing. We explained the superiority of LKCA from both convolution and attention perspectives, providing equivalent code implementations for each view. Experiments confirm that LKCA implemented from both the convolutional and attention perspectives exhibit equivalent performance. We extensively experimented with the LKCA variant of ViT in both classification and segmentation tasks. The experiments demonstrated that LKCA exhibits competitive performance in visual tasks. Our code has already been made publicly available at https://github.com/CatworldLee/LKCA.
Keywords:
@@
1 Introduction
Convolutional Neural Networks (CNNs) [1, 2] have been a common choice for visual encoders in modern visual models. However, the emergence of Vision Transformers (ViTs) [3] has expanded the options for the backbone network of models. ViTs have demonstrated excellent performance in various visual tasks [4, 5, 6, 7].
@@ 1 Introduction
CNN and ViT each possess unique advantages. CNN is fundamentally translation equivariance and shifting window invariance [8], providing robustness to subtle changes in object positions and orientations in images, leading to significant success in various computer vision tasks. In contrast, ViT effectively captures global context information through its self-attention mechanism [9], enabling the modeling of long-range dependencies [10] and contextual relationships [11]. Despite the significant progress made by CNN and ViT over the years, they still have some drawbacks and challenges. CNN may face difficulties in capturing remote contextual information [12], as its primary focus is on local feature extraction, potentially limiting its ability to understand global context [13]. This limitation can result in performance degradation in tasks that require global context information, such as scene understanding, image captioning, or fine-grained recognition. On the other hand, ViT lacks the inherent spatial inductive bias that CNN possesses [14], which is specifically designed to leverage the spatial locality of images.
In order to integrate the strengths of both CNN and ViT, we propose a Large Kernel Convolutional Attention (LKCA) mechanism. The implementation of LKCA is straightforward, replacing the self-attention operation with a single large kernel convolution operation. LKCA not only inherits the translational invariance and spatial inductive bias advantages of ConvNets but also leverages the global modeling and long-range dependency capabilities of ViT. Deconstructing LKCA from the perspective of attention can be considered as a novel approach of spatial attention. The approach closest to our work is Visual Attention Network (VAN) [15], where they replace the multi-head self-attention mechanism with three types of small kernel convolutions: DW-Conv, DW-DConv, and 1$\times$1-Conv. In contrast, our method is simpler, directly employing one large kernel convolution [12] to replace MHSA.
We extensively validated the effectiveness of LKCA in two visual tasks: image classification and semantic segmentation. In image classification, we conducted experiments on classic datasets including CIFAR10, CIFAR100, SVHN, and Tiny-ImageNet, demonstrating the efficacy of LKCA across these datasets. Notably, in the case of Tiny-ImageNet, we performed comparative experiments against prominent variants of Vision Transformers ViT and achieved compelling and competitive results. For semantic segmentation, we selected the ADE20K dataset and conducted ablation experiments under different framework configurations. Through comparisons with multiple backbone networks, we verified the superior performance of LKCA in terms of both parameter efficiency and accuracy. This further underscores the excellence of LKCA in addressing semantic segmentation tasks.
2 Related Work
2.1 Visual Attentions
@@
Attention mechanisms [16, 17, 18, 19, 20, 21] are inspired by the way humans naturally focus on important elements in their subconscious. Visual attention mechanisms [20, 21] can be viewed as a dynamic weighting process based on input image features. Visual attention methods can be broadly categorized into four types: channel attention [22, 23, 24, 25], spatial attention [26, 27, 28], temporal attention [29, 30], and branch attention [31, 32, 33], as well as other mixed combinations [34, 35, 36, 37, 38, 39, 40]. Non-Local Network [41] successfully employs self-attention for the first time to simulate non-local relationships in computer vision. CBAM [42] infers attention weights sequentially along both spatial and channel dimensions and can seamlessly integrate into any CNN architecture. ECA [22] module introduces a non-reductive local cross-channel interaction strategy, which efficiently enhances the attention module’s effectiveness through the effective implementation of 1D convolution. EncNet [23] significantly improves the spatial resolution of pixel-level semantic segmentation labels by introducing a context encoding module. OCNet [26] combines an interleaved sparse self-attention scheme with a traditional multi-scale context approach, effectively modeling dense relationships between pixels and providing richer contextual information. SAGAN [27] employs spatial attention to model long-range dependencies in image generation tasks. SCAN [29] achieves significantly superior performance in Video Person Re-Identification by refining both intra-sequence and inter-sequence feature representations through the introduction of temporal attention to video sequences. SKNet [31] enhances the model’s capability to capture objects at different scales by introducing branch attention to merge multiple branches with different kernel sizes.
2.2 Large Kernel ConvNets
@@
Compared to small-kernel CNNs [1, 43], large-kernel CNNs [44, 45, 46, 47, 48, 49, 50, 12, 15] possess a broader effective receptive field and stronger shape preferences. However, older models like Inception [44, 45, 46] fell out of favor after VGG-Net [47], leading to a decline in the popularity of large-kernel models. A notable research outcome is the Global Convolutional Network (GCN) [48], which employs very large convolutional kernels of 1$\times$×K and K×$\times$1 to optimize semantic segmentation tasks. However, there are reports suggesting that large kernels may compromise ImageNet performance. On the other hand, the Local Relation Network (LRNet) [49] introduces a spatial aggregation operator (LRLayer) as a replacement for standard convolution, which can be seen as a dynamic convolution. ConvMixer [50], on the contrary, uses convolutional kernels as large as 9$\times$9 to replace the ”mixer” components of ViT or MLP. Inspired by ViT, RepLKNet [12] utilizes a large 31$\times$31 convolutional kernel, and its outstanding performance is primarily attributed to the effective receptive fields constructed through large kernels [65]. It has been demonstrated that RepLKNet better exploits shape information compared to traditional CNNs. The Visual Attention Network (VAN) [15] employs three types of convolutions—DW-Conv, DW-DConv, and 1$\times$1 Conv—to expand the receptive field of convolutions and integrate them into the attention mechanism.
2.3 Large ERF ViTs
@@
The strength of ViTs lies in its large Effective Receptive Fields (ERFs). Transformer [51] was initially applied in the field of Natural Language Processing, and the first work to migrate its powerful representation capabilities to the Computer Vision domain was ViT [3]. Following the paradigm of ViT, a series of ViT variants [52, 53, 54, 55, 56, 57, 58, 59, 60, 61] have been proposed to enhance the performance of visual tasks. Visual Transformers excel at capturing long-range dependencies between patches, representing large receptive fields. However, ViT lacks the dependence on short-term relationships for 2D image local dependencies that CNNs inherently possess. Researchers have primarily focused on improving ViT’s modeling capability for local information. Transformer in Transformer (TNT) [52] enhances the feature representation capability of local patches through internal attention. Swin Transformer [53, 54] addresses local modeling challenges in adapting transformers from language to vision by introducing a hierarchical architecture with shift windowing. Twins [55] and CAT [56] alternately perform local and global attention layers layer-wise. Shuffle Transformer [57, 58] efficiently establishes inter-window connections through spatial shuffling with deep convolution and enhances short-term dependency modeling with neighboring window connections. RegionViT [59] introduces regions and local tokens with different patch sizes, enhancing local modeling through a region-to-local attention mechanism.
3 Method
3.1 Preliminary
3.1.1 Review of Large Kernel Convolution.
@@ 3.1 Preliminary
Ding et al. [12] revisit the paradigm of large kernel design in modern CNNs, drawing inspiration from the ViT. Ding et al. assert that employing a few large convolutional kernels instead of multiple smaller ones constitutes a more powerful design paradigm. Building upon this notion, the author introduces a novel CNN architecture named RepLKNet [12], featuring kernel sizes as large as 31$\times$31. In comparison to commonly used 3$\times$3 kernels, RepLKNet achieves performance results equivalent to or better than the Swin Transformer [53], with lower latency. The difference between small kernel convolution and large kernel convolution is shown in Figure 2
3.1.2 Comparision of LKC and MHSA.
@@
Convolution possesses a natural advantage wherein parameters are shared based on geographical location. Specifically, identical parameters of the same convolutional kernel act on every position within the receptive field during the sliding window process. This introduces equivariance and translation invariance to the convolutional operation. However, in Vision Transformer, the attention scores generated by the self-attention mechanism do not explicitly exhibit positional invariance. Attention weights are determined by the feature vectors themselves and the introduced positional encodings. While this enhances the model’s fitting capability, it significantly increases the complexity of training. Introducing a mechanism for sharing parameters in the attention mechanism of ViT may prove beneficial in alleviating training challenges.
@@
Figure 2: Illustration of the difference between small kernel convolution and large kernel convolution by constructing a 5x5 feature map, a 3x3 convolution kernel smaller than the feature map, and a 7x7 convolution kernel larger than the feature map. The distinction between small kernel convolution and large kernel convolution lies in the fact that, in small kernel convolution, all parameters of the kernel are involved in each correlation operation, while only a subset of feature map parameters participates in the computation. In the case of large kernel convolution, during each correlation operation, all parameters of the feature map are involved, and only a subset of the convolution kernel parameters participates in the computation.
3.2 Large Kernel Convolution Attention (LKCA)
3.2.1 Shared Weight Position Operation.
@@ 3.2 Large Kernel Convolution Attention (LKCA)
Firstly, we initialize a square matrix of size $N^{\frac{1}{2}}\times 2-1$ as learnable parameters, denoted as the shared parameter weights, where $N$ represents the number of patches, akin to the initialization of convolutional kernels. Subsequently, we initialize a window of size $N^{\frac{1}{2}}$ and traverse this window as it slides over the shared parameter weights matrix from the bottom-right corner to the top-right corner. At each sliding step, the matrix within the window is flattened row-wise into a row vector of size $(1,N)$ and stored. This process is iterated as the window slides, concatenating the row vectors obtained from each step. In total, the window slides $N$ times. Consequently, the concatenated matrix assumes a shape of $(N,N)$, aligning with the shape of the attention weight matrix. This concatenated matrix serves as the new attention score, undergoing matrix multiplication with the linearly mapped Value matrix. This step is referred to as the Shared Weight Matrix Attention. Listing 1.1 presents pseudocode for implementing LKCA from the attention perspective.
class Attention_ViT(nn.Module):
    def __init__(num_patches):
        super().__init__()
        n = num_patches
        H, W = n**0.5
        weight = nn.Parameter(torch.zeros(H*2-1, W*2-1))
    def forward(X):
        b, n, d = X.shape
        V = nn.Linear(X)
        for i in range(H):
            for j in range(W):
                start_row = H-1-i
                start_col = W-1-j
                window = weight[start_row:start_row+H, start_col:start_col+W]
                window_vector = window.flatten()
                attn_score[i*W+j, :] = window_vector
        Out = attn_score @ V
        return Out
Listing 1: Pseudocode from attention perspective
3.2.2 From the Perspectives of Convolution: Implementing LKCA.

### v3

@@ Abstract
Attention mechanisms have significantly advanced visual models by capturing global context effectively. However, their reliance on large-scale datasets and substantial computational resources poses challenges in data-scarce and resource-constrained scenarios. Moreover, traditional self-attention mechanisms lack inherent spatial inductive biases, making them suboptimal for modeling local features critical to tasks involving smaller datasets. In this work, we introduce Large Kernel Convolutional Attention (LKCA), a novel formulation that reinterprets attention operations as a single large-kernel convolution. This design unifies the strengths of convolutional architectures—locality and translation invariance—with the global context modeling capabilities of self-attention. By embedding these properties into a computationally efficient framework, LKCA addresses key limitations of traditional attention mechanisms. The proposed LKCA achieves competitive performance across various visual tasks, particularly in data-constrained settings. Experimental results on CIFAR-10, CIFAR-100, SVHN, and Tiny-ImageNet demonstrate its ability to excel in image classification, outperforming conventional attention mechanisms and vision transformers in compact model settings. These findings highlight the effectiveness of LKCA in bridging local and global feature modeling, offering a practical and robust solution for real-world applications with limited data and resources.
Index Terms:
@@
I Introduction
Vision Transformers (ViTs) have emerged as a powerful backbone network in modern visual models, introducing a paradigm shift from traditional convolutional architectures [1]. By leveraging self-attention mechanisms, ViTs excel at capturing long-range dependencies and global context, achieving state-of-the-art performance in resource-intensive tasks [2, 3, 4, 5, 6, 7, 8, 9, 10], such as large-scale image classification[2], high-resolution segmentation [3], and advanced feature learning [4]. However, their effectiveness often comes with substantial computational and data requirements, making them less ideal for data-constrained scenarios [4, 11].
@@ I Introduction
Conversely, Convolutional Neural Networks (CNNs) have remained a robust choice in environments with limited computational resources or data availability [12, 13]. Due to their strong inductive biases, such as spatial locality and translation invariance, CNNs are inherently efficient in learning from smaller datasets and exhibit robustness to subtle variations in object positions and orientations [14]. These properties have made CNNs a go-to solution for tasks in low-resource settings, including embedded vision applications and small-scale datasets. Given their complementary strengths, CNNs can be employed to enhance the performance of ViTs, particularly in data-scarce environments. For instance, CNNs can serve as efficient feature extractors to preprocess input data for ViTs, reducing the computational load on the transformer layers [15]. Additionally, hybrid architectures that integrate convolutional modules within ViTs have demonstrated promise in alleviating the limitations of ViTs, such as their lack of inherent spatial inductive bias [15], while retaining their ability to capture global context [16, 17]. This synergy enables the design of models that balance computational efficiency with high-performance capabilities, paving the way for their deployment across a wider range of visual tasks, from resource-constrained devices to large-scale systems.
In this work, we introduce a mechanism called Large Kernel Convolutional Attention (LKCA), designed to effectively combine the strengths of CNNs and ViTs on small datasets. Following the ViT structure, LKCA utilizes key techniques such as Patch Embedding, Positional Encoding, Multi-Head Self-Attention, Feedforward Networks in Transformer Encoder layers, and MLP Head for task output. To effectively incorporate the benefits of CNNs, we employ a simple yet highly efficient operation in LKCA: replacing the conventional self-attention operation in ViT with a single large-kernel convolution operation. This approach not only retains the advantages of translation invariance and spatial inductive bias in CNNs but also leverages the global modeling capability and long-range dependency handling of ViTs. With this large-kernel convolution, LKCA captures both local and global information more effectively, enhancing the model’s representational power. From the perspective of attention mechanisms, LKCA can be viewed as an innovative spatial attention method. It efficiently captures spatial features within local regions through the large-kernel convolution operation while also benefiting from the global modeling capabilities of ViT, thus significantly improving the model’s ability to understand and process complex structures in images. The approach most closely related to our work is the Visual Attention Network (VAN)[18], which replaces the multi-head self-attention mechanism with three types of small kernel convolutions: DW-Conv, DW-DConv, and 1$\times$1-Conv. By contrast, our method is simpler, employing a single large kernel convolution[19] to replace MHSA. We primarily benchmarked small and medium-sized models, as it was previously believed that ViT outperforms CNNs on large datasets and large models, while CNNs perform better than ViTs on ultra-small and specialized datasets. On the CIFAR-10 dataset, LKCA demonstrate superior improvements over the standard ViT across various experimental configurations.
More importantly, the LKCA design performs exceptionally well on downstream tasks. Firstly, in the image classification task, we conducted experiments on classic datasets, including CIFAR-10, CIFAR-100, SVHN, and Tiny-ImageNet, demonstrating the efficacy of LKCA across these datasets. Additionally, LKCA achieves remarkable results in semantic segmentation tasks, particularly with the UPerNet and FPN algorithms. On the ADE20K dataset, LKCA outperforms other models in terms of both aAcc and mIoU after 10,000 and 20,000 iterations. Comparisons with multiple backbone networks further verify the superior performance of LKCA in both parameter efficiency and accuracy. We attribute the high performance of LKCA to the complementary strengths of CNNs and ViTs. By combining CNN’s proficiency in local feature extraction with ViT’s ability to model global context, LKCA achieves exceptional performance across various visual tasks.
In this work, our main contributions are as follows:
  • We propose Large Kernel Convolutional Attention (LKCA), a novel mechanism that integrates large-kernel convolution into the ViT architecture to combine the local feature extraction advantages of CNNs with the global modeling capabilities of ViTs.
  • We demonstrate that LKCA achieves strong performance across various tasks, including image classification on datasets such as CIFAR-10, CIFAR-100, SVHN, and Tiny-ImageNet, as well as semantic segmentation on ADE20K, outperforming several existing backbone architectures in terms of accuracy and efficiency.
  • We address the limitations of both CNNs and ViTs by focusing on small and medium-sized models, offering an effective solution for data-limited and resource-constrained scenarios.
II Related Works
II-A Attention Mechanisms in Visual Models
@@
Attention mechanisms [20, 21, 22, 23, 24, 25, 26], originally inspired by the way humans subconsciously focus on and prioritize important aspects of their surroundings, have become a cornerstone in modern machine learning and deep learning research. These mechanisms mimic the human ability to allocate cognitive resources dynamically, making them highly effective in diverse tasks across Natural Language Processing (NLP), computer vision, and beyond. Visual attention mechanisms [24, 25], a critical subset of these techniques, can be conceptualized as a process of dynamically assigning weights to specific features within input images. This weighting mechanism helps models to focus on the most relevant parts of an image, significantly improving the interpretability and performance of various computer vision models. Visual attention methods are diverse and can be broadly classified into four primary categories: channel attention[27, 28, 29, 30], spatial attention[31, 32, 33], temporal attention[34, 35], and branch attention[36, 37, 38]. These categories often appear in conjunction with other innovative combinations [39, 40, 41, 42, 43, 44, 45], reflecting the increasing sophistication of attention-based methodologies. Non-Local Network [46] successfully employs self-attention for the first time to simulate non-local relationships in computer vision. CBAM [47] infers attention weights sequentially along both spatial and channel dimensions and can seamlessly integrate into any CNN architecture. ECA [27] module introduces a non-reductive local cross-channel interaction strategy, which efficiently enhances the attention module’s effectiveness through the effective implementation of 1D convolution. EncNet [28] significantly improves the spatial resolution of pixel-level semantic segmentation labels by introducing a context encoding module. OCNet [31] combines an interleaved sparse self-attention scheme with a traditional multi-scale context approach, effectively modeling dense relationships between pixels and providing richer contextual information. SAGAN [32] employs spatial attention to model long-range dependencies in image generation tasks. SCAN [34] achieves significantly superior performance in Video Person Re-Identification by refining both intra-sequence and inter-sequence feature representations through the introduction of temporal attention to video sequences. SKNet [36] enhances the model’s capability to capture objects at different scales by introducing branch attention to merge multiple branches with different kernel sizes.
II-B ConvNets With Large Kernel
@@
Compared to traditional small-kernel CNNs [12, 48], which have dominated deep learning models due to their efficiency and simplicity, large-kernel CNNs [49, 50, 51, 52, 53, 54, 55, 19, 18] offer a distinct advantage with their broader effective receptive field and stronger preferences for capturing global shape information. These characteristics make large-kernel models particularly appealing for tasks requiring extensive contextual understanding, such as segmentation and recognition. Despite these advantages, early large-kernel models, such as the Inception family [49, 50, 51], gradually lost favor following the success of architectures like VGG-Net [52], which demonstrated the efficiency of stacking smaller kernels. This shift in focus led to a temporary decline in the adoption of large-kernel designs in mainstream applications. A notable research outcome is the Global Convolutional Network (GCN) [53], which employs very large convolutional kernels of 1$\times$×K and K×$\times$1 to optimize semantic segmentation tasks. However, there are reports suggesting that large kernels may compromise ImageNet performance. On the other hand, the Local Relation Network (LRNet) [54] introduces a spatial aggregation operator (LRLayer) as a replacement for standard convolution, which can be seen as a dynamic convolution. ConvMixer [55], on the contrary, uses convolutional kernels as large as 9$\times$9 to replace the ”mixer” components of ViT or MLP. Inspired by ViT, RepLKNet [19] utilizes a large 31$\times$31 convolutional kernel, and its outstanding performance is primarily attributed to the effective receptive fields constructed through large kernels. It has been demonstrated that RepLKNet better exploits shape information compared to traditional CNNs. The Visual Attention Network (VAN) [18] employs three types of convolutions—DW-Conv, DW-DConv, and 1$\times$1 Conv—to expand the receptive field of convolutions and integrate them into the attention mechanism.
II-C ViTs with Large ERF
@@
The strength of ViTs lies in their ability to construct large Effective Receptive Fields (ERFs), enabling them to capture global dependencies effectively. Initially, the Transformer architecture [56] was developed for NLP, revolutionizing the way sequential data was modeled through its self-attention mechanism. Its migration to the domain of Computer Vision began with ViT [1], which demonstrated the powerful representation capabilities of transformers in handling image data. ViT divides images into patches and treats them as sequences, allowing the transformer architecture to process images in a manner analogous to text in NLP. Building on this foundation, numerous ViT variants [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] have been proposed to enhance performance on a wide range of visual tasks. These variants have introduced various improvements aimed at addressing the unique challenges posed by image data, particularly the need to model both global and local dependencies effectively. The global modeling capabilities of ViTs make them exceptionally well-suited for capturing long-range relationships between image patches, effectively creating large receptive fields. However, the lack of inherent mechanisms for modeling short-term relationships, as found in CNNs, has been a notable limitation of ViTs. CNNs excel at capturing local dependencies through their spatially constrained convolutional operations, a property that ViTs inherently lack due to their patch-based processing. To address these challenges, researchers have focused on enhancing ViT’s ability to model local information, leading to innovative designs and architectures. Transformer in Transformer (TNT) [57] enhances the feature representation capability of local patches through internal attention. Swin Transformer [58, 59] addresses local modeling challenges in adapting transformers from language to vision by introducing a hierarchical architecture with shift windowing. Twins [60] and CAT [61] alternately perform local and global attention layers layer-wise. Shuffle Transformer [62, 63] efficiently establishes inter-window connections through spatial shuffling with deep convolution and enhances short-term dependency modeling with neighboring window connections. RegionViT [64] introduces regions and local tokens with different patch sizes, enhancing local modeling through a region-to-local attention mechanism.
III Approach
III-A Preliminary
III-A1 Review of Large Kernel Convolution
@@
Ding et al. [19] revisit the paradigm of large kernel design in modern CNNs, drawing inspiration from the ViT. Ding et al. assert that employing a few large convolutional kernels instead of multiple smaller ones constitutes a more powerful design paradigm. Building upon this notion, the author introduces a novel CNN architecture named RepLKNet [19], featuring kernel sizes as large as 31$\times$31. In comparison to commonly used 3$\times$3 kernels, RepLKNet achieves performance results equivalent to or better than the Swin Transformer [58], with lower latency. The difference between small kernel convolution and large kernel convolution is shown in Figure 2.
III-A2 Comparision of LKC and MHSA
@@
Convolution possesses a natural advantage wherein parameters are shared based on geographical location. Specifically, identical parameters of the same convolutional kernel act on every position within the receptive field during the sliding window process. This introduces equivariance and translation invariance to the convolutional operation. However, in Vision Transformer, the attention scores generated by the self-attention mechanism do not explicitly exhibit positional invariance. Attention weights are determined by the feature vectors themselves and the introduced positional encodings. While this enhances the model’s fitting capability, it significantly increases the complexity of training. Introducing a mechanism for sharing parameters in the attention mechanism of ViT may bring the expected benefits but without increasing optimization complexity.
@@
Figure 2: Participation differences in kernel convolutions. Illustration of the difference between small kernel convolution and large kernel convolution by constructing a 5x5 feature map, a 3x3 convolution kernel smaller than the feature map, and a 7x7 convolution kernel larger than the feature map. The distinction between small kernel convolution and large kernel convolution lies in the fact that, in small kernel convolution, all parameters of the kernel are involved in each correlation operation, while only a subset of feature map parameters participates in the computation. In the case of large kernel convolution, during each correlation operation, all parameters of the feature map are involved, and only a subset of the convolution kernel parameters participates in the computation.
III-B Large Kernel Convolutional Attention (LKCA)
III-B1 Shared Weight Position Operation
@@
Firstly, we initialize a square matrix of size $N^{\frac{1}{2}}\times 2-1$ as learnable parameters, denoted as the shared parameter weights, where $N$ represents the number of patches, akin to the initialization of convolutional kernels. Then, we initialize a window of size $N^{\frac{1}{2}}$ and traverse this window as it slides over the shared parameter weights matrix from the bottom-right corner to the top-right corner. At each sliding step, the matrix within the window is flattened row-wise into a row vector of size $(1,N)$ and stored. This process is iterated as the window slides, concatenating the row vectors obtained from each step. In total, the window slides $N$ times. Consequently, the concatenated matrix assumes a shape of $(N,N)$, aligning with the shape of the attention weight matrix. This concatenated matrix serves as the new attention score, undergoing matrix multiplication with the linearly mapped Value matrix. This step is referred to as the Shared Weight Matrix Attention. Algorithm 1 presents pseudocode for implementing LKCA from the attention perspective.
Algorithm 1 Forward propagation process of Attention-ViT
1:function Attention_ViT_Forward($X$)
2:     input: $X$ – input tokens, where $X\in\mathbb{R}^{b\times n\times d}$
3:     initialize: $H,W\leftarrow\sqrt{n}$, $weight\leftarrow$ zero tensor of shape $(H*2-1,W*2-1)$
4:     $V\leftarrow\text{Linear}(X)$
5:     for $i=0$ to $H-1$ do
6:         for $j=0$ to $W-1$ do
7:              $start\_row\leftarrow H-1-i$
8:              $start\_col\leftarrow W-1-j$
9:              $window\leftarrow weight[start\_row:start\_row+H,start\_col:start\_col+W]$
10:              $window\_vector\leftarrow\text{flatten}(window)$
11:              $attn\_score[i*W+j,:]\leftarrow window\_vector$
12:         end for
13:     end for
14:     $Out\leftarrow attn\_score@V$
15:     return $Out$
16:end function
III-B2 From the Perspective of Convolution: Implementing LKCA